# Attosecond physics hidden in Cherenkov radiation

### v1

@@ Introduction
Cherenkov radiation (ChR) by charged particles in media [1, 2, 3, 4, 5, 6, 7, 8, 9] is the simplest example of a wide range of phenomena embracing transition, diffraction, Smith-Purcell, and other mechanisms of photon emission [10, 11, 12, 13]. Their common microscopic origin is atomic bremsstrahlung [14] due to dynamic polarization of atomic shells by the field of the charge, and the characteristic time scales are femto- and attoseconds, typical for the AC Stark effect [15]. Despite the growing interest to quantum features of ChR [16, 17, 18], a solid link between microscopic and macroscopic theories remains elusive, and the atomic dynamics stays hidden even in the fully quantum treatment of ChR in momentum space and of its generalizations.
Here we point out that one can access atomic time scales in ChR by using a quantum theory in phase space. Similar to quantum optics [19], we employ a Wigner function to characterize the emitted photon, which depends on a momentum ${\bf k}$, on a point ${\bf r}$ in space and a time instant $t$. We demonstrate how to probe the field in the pre-wave (or formation) zone [20, 11] in which the partial waves interfere, spread, and the Cherenkov cone is not formed yet. Most strikingly, our theory establishes a link between the evolving coherence length of the electron packet, the Cherenkov flash duration, and a quantum temporal delay that the photon experiences in medium. This delay can be either positive or negative and belongs to the attosecond range, typical for atomic excitation processes studied by the attosecond spectroscopy and metrology [21]. The quantum shift in the photon arrival time, coupled with the finite flash duration and the spreading dynamics, unveils intricate coherence properties that enable refined temporal control in quantum emission processes, enriching the landscape of quantum optics, ultrafast physics, and of applications of ChR and related phenomena. The system of units $\hbar=c=1$ is used, the electron velocity is $\beta=u_{p}/c\equiv u_{p}$, $m=0.511$ MeV is the electron mass, $1/m\equiv\hbar/mc=3.86\times 10^{-11}$ cm, and $t_{c}=1/m\equiv\hbar/mc^{2}\approx 1.3\times 10^{-21}\,\text{s}$.
@@
Here $|S_{fi}^{(\text{pw})}|^{2}$ defines the emission rate in momentum space where the phase $\zeta_{fi}$ does not contribute to the probability, although it is non-vanishing even in the lowest order of the perturbation theory (Sec.5 in Supplementary). The result of the calculations is (Sec.3 in Supplementary)
@@
$\sigma_{t}(0)\sim 10\,\text{as}-1\,\text{fs}.$ (18)
@@
In practice, achieving attosecond flash durations requires electrons generating radiation with no spreading before the target, just after the emission from a cathode. A vacuum gap of even a few millimeters after the source would yield micrometer-sized electron packets entering the target, which results in picosecond durations (subpicosecond with acceleration to a few MeV) measurable by streak cameras. State-of-the-art Cherenkov counters at accelerators have picosecond time resolution [40, 42], the subfemtosecond resolution can be achieved at X-ray free-electron lasers [43, 44] or with the frequency-resolved optical gating [45], whereas it is nanoseconds for Cherenkov telescopes employed in gamma-ray astronomy [46] where spreading effects are notable. Attosecond photon pulses – including twisted photons with orbital angular momentum – are usually obtained through high-harmonic generation in the extreme-ultraviolet and even soft X-ray ranges [47, 48], enabling vortex electrons generation via photoionization [48]. The ChR can be a source of twisted photons [17, 49], also in the soft X-ray range, and shaping the spatial and temporal profiles of the electron wave function offers refined tuning of phase-space profile of the photons.
Close to the Cherenkov angle, the flash duration is directly tied to the electron coherence length upon entering the medium, opening avenues for contolling this duration by selecting packets of desired length and accelerating them in the low-current regime with no space-charge effects, typical for TEMs. Alternatively, measuring photon pulse durations can provide a novel method for determining electron packet length, a complementary approach to interferometry that is reminiscent of bunch length measurements via classical coherence [13]. Naturally, ChR serves as an exemplary case within a broader class of media-induced emission phenomena where attosecond time scales are accessible via phase-space analysis.
@@
Author contributions. All authors contributed extensively to this work. Dm.K. conceived the idea, developed the theoretical framework, wrote the manuscript with input from all the authors and acquired the funding. D.G., G.S., A.C., and D.K. checked the calculations and interpreted the results in detail. D.G., A.C., and G.S. assisted with the figures, while A.C., Dm.K., and G.S. refined the presentation of results and theoretical details.

### v2

@@ Introduction
Cherenkov radiation (ChR) by charged particles in media [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] is the simplest example of a wide range of phenomena embracing transition, diffraction, Smith-Purcell radiation, and other mechanisms of photon emission [11, 12, 13, 14]. Their common microscopic origin is atomic bremsstrahlung [15] due to dynamic polarization of atomic shells by the field of the charge, and the characteristic time scales are femto- and attoseconds, typical for the AC Stark effect [16]. Along with numerous applications of classical ChR in particle detection [17], neutrino telescopes [18, 19], in gamma-ray astronomy [20, 21], and other fields, it has also recently attracted attention of biomedical community as a new tool for molecular imaging and therapy in cancer treatment [22, 23, 24]. However, despite the growing interest to inherently quantum features of ChR [25, 26, 27], a solid link between microscopic and macroscopic theories remains elusive, and the atomic dynamics stays hidden even in the fully quantum treatment of ChR in momentum space.
Here we point out that one can access atomic time scales in ChR and in its generalizations by using a quantum theory in phase space where – similar to quantum optics [28] – we employ a Wigner function to characterize the emitted photon field. We demonstrate how to probe the field in the pre-wave (or formation) zone [29, 12, 30] in which the partial waves interfere, spread, and the Cherenkov cone is not formed yet. Most strikingly, our theory establishes a link between the evolving coherence length of the electron packet, the Cherenkov flash duration, and a quantum temporal delay that the photon experiences in medium. This delay can be either positive or negative and belongs to the attosecond range, typical for atomic excitation processes studied by the attosecond spectroscopy and metrology [31]. The quantum shift in the photon arrival time, coupled with the finite flash duration and the spreading dynamics, unveils intricate coherence properties that enable refined temporal control in quantum emission processes, enriching the landscape of quantum optics, ultrafast physics, and of applications of ChR and related phenomena. The system of units $\hbar=c=1$ is used, the electron velocity is $\beta=u_{p}/c\equiv u_{p}$, $m=0.511$ MeV is the electron mass, $1/m\equiv\hbar/mc=3.86\times 10^{-11}$ cm, and $t_{c}=1/m\equiv\hbar/mc^{2}\approx 1.3\times 10^{-21}\,\text{s}$.
@@
Here $|S_{fi}^{(\text{pw})}|^{2}$ defines the emission rate in momentum space where the phase [34, 35, 36] $\zeta_{fi}$ does not contribute to the probability, although it is non-vanishing even in the lowest order of the perturbation theory (Sec.5 in Supplementary). The result of the calculations is (Sec.3 in Supplementary)
@@
$\displaystyle\sigma_{t}(0)\sim 10\,\text{as}-1\,\text{fs}.$ (18)
@@
In practice, achieving attosecond flash durations requires electrons generating radiation with no spreading before the target, just after the emission from a cathode. A vacuum gap of even a few millimeters after the source would yield micrometer-sized electron packets entering the target, which results in picosecond durations (subpicosecond with acceleration to a few MeV) measurable by streak cameras. State-of-the-art Cherenkov counters at accelerators have picosecond time resolution [53, 55], the subfemtosecond resolution can be achieved at X-ray free-electron lasers [56, 57] or with the frequency-resolved optical gating [58], whereas it is nanoseconds for Cherenkov telescopes in gamma-ray astronomy [21] where spreading effects are notable. Attosecond photon pulses – including twisted photons with orbital angular momentum – are usually obtained through high-harmonic generation in the extreme-ultraviolet and even soft X-ray ranges [59, 60], enabling vortex electrons generation via photoionization [60]. The ChR can be a source of twisted photons [26, 61], also in the soft X-ray range, and shaping the spatial and temporal profiles of the electron wave function offers refined tuning of phase-space profile of the photons.
Close to the Cherenkov angle, the flash duration is directly tied to the electron coherence length upon entering the medium, opening avenues for contolling this duration by selecting packets of desired length and accelerating them in the low-current regime with no space-charge effects, typical for TEMs. Alternatively, measuring photon pulse durations can provide a novel method for determining electron packet length, a complementary approach to interferometry that is reminiscent of bunch length measurements via classical coherence [14]. The developed framework provides visualization of the emitted photon energy in real space and time, which is particularly desirable for biomedical applications of ChR in radiotherapy [23, 24] where only the classical Tamm-Frank theory in the far field has been used so far, apparently overestimating the process intensity in the formation zone. Naturally, ChR serves as an exemplary case within a broader class of media-induced emission phenomena where attosecond time scales are accessible via the phase-space analysis.